Suppose you are given the side lengths of a right angle triangle, which length is the length of the hypotenuse. How do you know?


1
Expert's answer
2022-04-08T14:01:15-0400

Pythagorean theorem:

"a^2+b^2=c^2,"


where a, b - side lengths;

c - length of hypotenuse.


"c=\sqrt{a^2+b^2}."